You hate when rabbits run into your yard and eat your garden. To prevent this, you can rent a ferocious wolf to scare the rabbits out of your yard! At Wolf "Em Up, they will deliver a live wolf to your door for 17.50 and you can keep the lively fellow for 2.25 per day. At Fangs R Us they will deliver a vicious pal for you for 24.25 and you can keep him for just 1.50 a day. ( At both businesses they will prorate a rental fee for partial days.
a. On what day number will be number of days rented and the cost of renting a wolf be at both businesses
b. What will the be on that day?
I am confused on how to start this problem I know that I must compare the businesses but I do not know where to start. I am a step by step person.

Answer by JulietG(1812)

The base cost for company A is $17.50 with a daily cost of $2.25
The base cost for company B is $24.25 with a daily cost of $1.50
A = 17.5 + 2.25x (where x is the number of days)
B = 24.25 + 1.50x
You want to know when they'll be equal
17.5 + 2.25x = 24.25 + 1.5x
Subtract 1.5x from each side
17.5 + .75x = 24.25
Subtract 17.5 from each side
.75x = 6.75
Divide each side by .75
x = 9
On the 9th day, the cost will be the same.
Company A = 17.5 + (9*2.25) = $37.75
Company B = 24.25 + (9*1.50) = $37.75
Fewer than 9 days and Company A is the better deal.
More than 9 days and Company B is the better deal.